OpenStudy (anonymous):

If t varies as v, and t = 2 4/7 when v = 13/14 , find v when t = 2 1/4. v= direct variation? PLEASE HELP!

OpenStudy (anonymous):

t v 24/7 13/14 21/4 hence v={(13/14)/(24/7)}*(21/4) v=19/64

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

didnt i show you this yesterday???

OpenStudy (anonymous):

nice pic btw...I don't know @sarkar...I don't hink so bc it's a new lesson